OpenStudy (anonymous):

Solve the exponential equation. Express the solution in terms of logarithms. 7^(4x+2)=6^(x-2)

OpenStudy (anonymous):

\[\ln (7^(4x+2))=\ln (6^(x-2))\] so \[(4x+2)*\ln(7)=(x-2)\ln(6)\] can you finish it by yourself ?!

OpenStudy (anonymous):

no im a little confused

OpenStudy (anonymous):

wait, do you put all x's on the left?

OpenStudy (anonymous):

yeaaaah ! you have a problem with that ?! :)

OpenStudy (anonymous):

do you have to divide by x-2 on both sides?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ln(a^b)=bln(a) and consider that ln(7) and ln(6) two real numbers...you can solve equations like : "4x+6=-3x+2" put x's on the left and you will get the answer !

OpenStudy (anonymous):

which part you didn't understand ?! :)

OpenStudy (anonymous):

I didnt understand how to get the final answer ;)